Assume that there is no storage or other carrying costs/value for corn. Assume that corn is not perishable, that is it can be stored and used later and that the continuous compounded risk-free rate is 20% per year.

If a 6-month corn future currently costs $4.75 a bushel and the spot price currently costs $4.50.

Suppose that investors had to pay a short sale fee whenever they short a contract. Assume that this fee was calculated like a carrying cost.

So that if you shorted corn you would have to pay short-selling costs when you enter the shorting contract.

Find the value of short selling costs that is consistent with no arbitrage opportunities.
Answer #1

Risk Free Rate = 20% compounded continuously

r = 0.2

t = 6/12 = 0.5

Spot = $ 4.50

6 months Future = $ 4.75

Future price = Spot × ert - Short selling cost

4.75 = 4.5 × e0.2×0.5 - Short selling cost

Short selling cost = 4.5 × 1.1052 - 4.75

Short selling cost = $ 0.2233